OpenStudy (loser66):

Find general formula for \(a_{n+2}=9a_{n+1}-8a_n\) where \(a_0=1, a_1=8\) Please, help

OpenStudy (loser66):

I got a2=64 a3= 568 a4=4600 but still not get the formula for \(a_n\) yet,

OpenStudy (loser66):

@ikram002p

OpenStudy (rational):

you want to use generating functions ?

OpenStudy (loser66):

I want recursive formula for an

OpenStudy (loser66):

\(a_n = 8^n\) is correct for n =0,1,2 only

OpenStudy (loser66):

it changes when n =3 and up

OpenStudy (loser66):

and I have \(a_3 = 73a_1-72a_0\) \(a_4= 585a_1-584a_0\)

OpenStudy (loser66):

but need general form for an

OpenStudy (loser66):

brb

OpenStudy (ikram002p):

;_; Sorry Hoa can't help now :(

OpenStudy (rational):

a_0 = 1 a_1 = 8 a_2 = 9*8 - 8*1 = 64 a_3 = 9*64 - 8*8 = 512 a_4 = 9*512 - 8*64 = 4096

OpenStudy (rational):

\(a_n = 8^n\) is a correct formula